OpenStudy (anonymous):

how long will it take an object to fall freely from a bridge 1053ft above the water using the formula s=16t^2

OpenStudy (anonymous):

what does "s" represent here? :)

OpenStudy (anonymous):

it is just part of the formula that i am given would it represent the 1053 ft?

OpenStudy (anonymous):

could be, if so then you can find t, that's why I have asked lol :)

OpenStudy (anonymous):

find t how?

OpenStudy (anonymous):

can you set it up fo rme?

OpenStudy (anonymous):

if s = represents the height in this equation then solve for t:\[s = 16t^2 \rightarrow 1053 = 16t^2\]\[t^2 = \frac{1053}{16} \rightarrow t = \sqrt(\frac{1053}{16}) \approx 8.11 \secs\] that if S = height in the equation ^_^

OpenStudy (anonymous):

ok, thank you very much.